Solve the rational equation
x + 1 = 5
_ _
x 2
and check

Answer by checkley77(12569)

x + 1 = 5
_ _
x 2
(x*x+1)/x=5/2
(x^2+1)/2=5/2
x^2+1=5
x^2+1-5=0
x^2-4=0
(x-2)^2=0
x-2=0
x=2 ans.
Proof:
2+1/2=5/2
(2*2+1)/2=5/2
(4+1)/2=5/2
5/2=5/2